Clark Zealand Co. has total fixed costs of $160,000. Its product sells for $40 per unit and variable costs amount to $30 per unit. What is the break-even point in dollar sales? Use the formula(s) that were introduced in this chapter. Hint: the BV point in dollar sales will be a number greater than $500,000
Answer #1

Contribution margin=Sales-Variable costs

=(40-30)=$10

Contribution margin ratio=Contribution margin/Sales

=(10/40)=0.25

Hence breakeven sales=Fixed costs/Contribution margin ratio

(160,000/0.25)

which is equal to

=$640,000.